The Foundational 15 [L09-1, LO9-2, L09-4, LO9-5, L09-6) [The following information applies to the questions displayed below.)12. What amounts of advertising, sales salaries and commissions, and shipping expenses would be included in the companys fle13. What is the spending variance related to advertising? (Indicate the effect of each variance by selecting F for favorabl14. What is the spending variance related to sales salaries and commissions? (Indicate the effect of each variance by selecti15. What is the spending variance related to shipping expenses? (Indicate the effect of each variance by selecting F for fa
Answer #1

12) Amount in Flexible budget

Advertising 200000
Sales salaries and commission (100000+30000*12) 460000
Shipping expense (30000*3) 90000

13) Spending variance for advertising = 200000-210000 = 10000 U

14) Spending variance for sales salaries and commission = 460000-455000 = 5000 F

15) Spending variance for shipping expense = 90000-115000 = 25000 U